OpenStudy (anonymous):

a marble is drawn at random from a box that contains 8 red, 15 blue, and 7 white marbles. whats the probability of getting a white a red marble?

OpenStudy (anonymous):

8+7/8+7+15=15/30=1/2

OpenStudy (anonymous):

ment red marble sorry!

OpenStudy (anonymous):

oh ok so that is 8/7+8+15=8/30=4/15.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome.